OpenStudy (anonymous):

-17-1+27-39=-2-2m-8+3m

OpenStudy (anonymous):

-30=-10+m -20=m 0=20+m

OpenStudy (anonymous):

-17-1+27-39=-2-2m-8+3m -18-12=-10+m -30=-10+m -20=m

OpenStudy (anonymous):

-17-1+27-39 = -30 -2-2m-8+3m = -10 +m -> m = -30 + 10 = -20